The frequency of the light emitted by a galaxy receding from the earth is measured
to be 1.5×103 MHz. Assuming that the wavelength of the light source is 22.5 cm,
calculate how fast the galaxy is receding from the earth?
1
Expert's answer
2020-04-02T04:57:47-0400

We are given

"\nu=1.5 \cdot 10^{3} MHz"

"\lambda_0=0.225 m"

Define the radiation frequency of a stationary galaxy

"\nu_0=\frac{c}{\lambda_0}=\frac{3 \cdot 10^{8}m/s}{0.225 m}=1.333 \cdot 10^{3}MHz"

Using the formula for the Doppler effect

"\nu=\frac{\nu_0}{1-\frac{V_r}{c}}"

then we will find "V_r"

"V_r=\frac{\nu-\nu_0}{\nu}\cdot c"

"V_r=\frac{1.5 \cdot 10^{3}-1.333 \cdot 10^{3}}{1.5 \cdot 10^{3}}\cdot 3 \cdot 10^{8}=3.34 \cdot 10^{7}m/s=0.111 \cdot c"